# revtex4-2Repair the float

### v1

@@ III.2 Coexistent magnetization and Neel structures
Let us go back to link correlations for the strained case and take a closer look at Fig. 4. Note that for angles where $\zeta$ or $E_{0}$ have a minimum (right panels of Fig. 4), the formation of 1D correlated lines conspire to leave the sites between lines only loosely coupled to their neighbors. This suggests the coexistence of antiferromagnetic links and paramagnetic sites in the lattice. One would further expect these two types of sites (free and correlated) to respond differently to an applied magnetic field. The (antiferromagnetically) correlated spins would require a larger field to be polarized, as the field should compete with the effective exchange interaction. On the other hand, the loose spins would align readily with the field.
Figure 7: Magnetization $M^{z}_{j}$ for $V_{z}=0.05$ and $\theta=0$ (a), $\theta={\pi}/{6}$ (b), $\theta={\pi}/{3}$ (c), $\theta={\pi}/{2}$ (d), $\theta={2\pi}/{3}$ (e) and $\theta={5\pi}/{6}$ (f). For all panels, we set $U=5$ and $\epsilon=0.30$. Faint dots on the right panels (where $E_{0}$ has minima in Fig. 3) correspond to sites belonging to Neel-like lines, while dark red dots correspond to loose sites readily polarized by the field.
Figure 8: Local magnetization $M^{z}_{j}$ for $V_{z}=0.05$ and strain applied along $\theta=\pi/3$ (left) and $\theta=\pi/6$ (right). Upper and lower panels show results for $U=10$ and $U=20$, respectively. As in Fig. 7, faint color dots on the right panels correspond to sites belonging to Neel AF lines, while dark red dots correspond to loose sites more easily polarized by the field. For all pannels, $\epsilon=0.30$.
@@
To verify this reasoning, we apply a small magnetic field along the $z$-direction to the entire system and analyze the behavior of the spatially-resolved magnetization over the lattice. Figure 7 shows the local $z$-magnetization $M_{j}^{z}=\langle S_{j}^{z}\rangle$ for the ground state in a small applied field $V_{z}=0.05$ and for the same strain angles as in Fig. 4. Colorful dots at each site of the lattice indicate whether that spin is polarized towards $+\hat{\boldsymbol{z}}$ (red), $-\hat{\boldsymbol{z}}$ (blue) or unpolarized (white), as per the color scale. Let us analyze what is shown in Fig. 7. Starting with the case of $\theta=0$ [Fig. 7(a)] we observe a pattern consisting of alternating positive/negative $M_{j}^{z}$ for $j$ on the line along the vectors ${\boldsymbol{a}}_{1}$ or ${\boldsymbol{a}}_{2}$ with all positive $M_{j}^{z}$ along ${\boldsymbol{a}}_{3}$. This pattern can be connected with Fig. 4(a) as follows: since the weakest link correlations are seen there along ${\boldsymbol{a}}_{3}$, the application of a magnetic field easily polarizes all spins parallel to the field for sites along ${\boldsymbol{a}}_{3}$; in contrast, sites along ${\boldsymbol{a}}_{1}$ and ${\boldsymbol{a}}_{2}$ align antiferromagnetically to those polarized by the field. The AF links remain intact because for this weak field, it is still energetically favorable to form the Neel lines along with strong link correlations. Similar conclusions can be made for the cases of $\theta=\pi/3$ and $\theta=2\pi/3$, except that the weakest links in these cases lie along ${\boldsymbol{a}}_{1}$ and ${\boldsymbol{a}}_{2}$, respectively. We emphasize that the sites aligned with the field and lying along ${\boldsymbol{a}}_{3}$ in Fig. 7(a) are consistent with the formation of AF lines along ${\boldsymbol{a}}_{1}$ and ${\boldsymbol{a}}_{2}$ in Fig. 4(a), and that the AF many-body correlations are not suppressed for small fields, as one would expect.
@@
The coexistence of distinct magnetic structures, strongly correlated links and easily polarizable loose sites is expected to be more pronounced for larger values of $U$. To test this, Fig. 8 shows the magnetization $M_{j}^{z}$ for $U=10$ (upper row) and $U=20$ (lower row) for the same magnetic field $V_{z}=0.05$. Left and right panels correspond to $\theta=\pi/3$ and $\theta=\pi/6$, respectively, corresponding to the lattice strains in Fig. 5. We again observe on the left panels that the weakly correlated sites along the $\theta$ direction (${\boldsymbol{a}}_{1}$) become easily polarized by the field, while the neighboring sites are/remain strongly AF correlated along ${\boldsymbol{a}}_{2}$ and ${\boldsymbol{a}}_{3}$ [as in Fig. 7(c)]. The right panels in Fig. 8 show that the energy minimizes on alternating diagonal lines that weakly polarize in the field while coexisting with strongly polarized sites through the structure, similarly to Fig. 7(b).
One may wonder whether the applied magnetic field destroys the many-body spin correlations. We expect the correlations to be robust along the well defined Neel lines. To confirm this, Fig. 9 shows the link correlations in the presence of a field for systems as in Fig. 4(a) and (b). Comparing these two figures, it is clear that correlations along the strong links are indeed preserved for the applied field $V_{z}=0.05$. Interestingly, the correlations along the lower edge are affected more, likely due to their fewer neighbors. Notice again that the Neel lines are parallel to the lattice vector ${\boldsymbol{a}}_{2}$ in Fig. 9(b) and 4(b) as the sites in between adjacent lines are weakly correlated to their neighbors and are then strongly polarizable by the field, as seen in Fig. 7(b). These sites behave as a collection of paramagnetic (noninteracting) polarizable spins, while the sites belonging to Neel lines form a separate set of AF 1D chains. The coexistence of these two classes of sites and the dependence of their appearance and orientation with applied strain represent the main unexpected results of this work.
@@ IV Conclusions
We have studied a kagome nanoribbon using a Hubbard model to describe how anisotropies induced by uniaxial strain break the symmetry and affect the correlations due to electronic interactions. We find that even for relatively weak interactions, the competing correlations and frustration develop coexisting systems of strongly antiferromagnetic correlations between neighboring sites and weakly connected sites that are easily polarizable in a magnetic field. The spatial distribution of the strong links and interspersed paramagnetic sites depends strongly on the orientation of the uniaxial strain, and opens the possibility of exploring an experimentally tunable probe that can give rise to different correlated behavior in a given lattice. The need to consider charge fluctuations to define a function that quantifies correlations, especially for weak $U$ values, and that complements $f^{G}$, remains an interesting theoretical question. With the growing interest in metallic kagome materials displaying strongly correlated behavior, it may be possible to apply external strain fields and explore some of the phenomena our models describe. It would be interesting, moreover, to explore how different Fermi levels and proximity to van Hove singularities in different materials affect this behavior.
@@ Appendix A Strained couplings
Now, by defining $\alpha_{\pm}\equiv\left(1\pm\nu\right)/2$, we can recast Eq. (7) as

### v2

@@ III.2 Coexistent magnetization and Neel structures
Let us go back to link correlations for the strained case and take a closer look at Fig. 4. Note that for angles where $\zeta$ or $E_{0}$ have a minimum (right panels of Fig. 4), the formation of 1D correlated lines conspire to leave the sites between lines only loosely coupled to their neighbors. This suggests the coexistence of antiferromagnetic links and paramagnetic sites in the lattice. One would further expect these two types of sites (free and correlated) would respond differently to an applied magnetic field. The (antiferromagnetically) correlated spins would require a larger field to be polarized, as the field should compete with the effective exchange interaction. On the other hand, the loose spins would align readily with the field.
@@
To verify this reasoning, we apply a small magnetic field along the $z$-direction to the entire system and analyze the behavior of the spatially-resolved magnetization over the lattice. Figure 7 shows the local $z$-magnetization $M_{j}^{z}=\langle S_{j}^{z}\rangle$ for the ground state in a small applied field $V_{z}=0.05$ and for the same strain angles as in Fig. 4. Colorful dots at each site of the lattice indicate whether that spin is polarized towards $+\hat{\boldsymbol{z}}$ (red), $-\hat{\boldsymbol{z}}$ (blue) or unpolarized (white), as per the color scale. Let us analyze what is shown in Fig. 7. Starting with the case of $\theta=0$ [Fig. 7(a)] we observe a pattern consisting of alternating positive/negative $M_{j}^{z}$ for $j$ on the line along the vectors ${\boldsymbol{a}}_{1}$ or ${\boldsymbol{a}}_{2}$ with all positive $M_{j}^{z}$ along ${\boldsymbol{a}}_{3}$. This pattern can be connected with Fig. 4(a) as follows: since the weakest link correlations are seen there along ${\boldsymbol{a}}_{3}$, the application of a magnetic field easily polarizes all spins parallel to the field for sites along ${\boldsymbol{a}}_{3}$; in contrast, sites along ${\boldsymbol{a}}_{1}$ and ${\boldsymbol{a}}_{2}$ align antiferromagnetically to those polarized by the field. The AF links remain intact because, for this weak field, it is still energetically favorable to form the Neel lines along with strong link correlations. Similar conclusions can be made for the cases of $\theta=\pi/3$ and $\theta=2\pi/3$, except that the weakest links in these cases lie along ${\boldsymbol{a}}_{1}$ and ${\boldsymbol{a}}_{2}$, respectively. We emphasize that sites aligned with the field and lying along ${\boldsymbol{a}}_{3}$ in Fig. 7(a) are consistent with the formation of AF lines along ${\boldsymbol{a}}_{1}$ and ${\boldsymbol{a}}_{2}$ in Fig. 4(a), and that the AF many-body correlations are not suppressed for small fields, as one would expect.
@@
The coexistence of distinct magnetic structures, strongly correlated links and easily polarizable loose sites is expected to be more pronounced for larger values of $U$. To test this, Fig. 8 shows $M_{j}^{z}$ for $U=10$ (upper row) and $U=20$ (lower row), for the same magnetic field $V_{z}=0.05$. Left and right panels correspond to $\theta=\pi/3$ and $\theta=\pi/6$, respectively, corresponding to the lattice strains in Fig. 5. We again observe on the left panels that the weakly correlated sites along the $\theta$ direction (${\boldsymbol{a}}_{1}$) become easily polarized by the field, while the neighboring sites are/remain strongly AF correlated along ${\boldsymbol{a}}_{2}$ and ${\boldsymbol{a}}_{3}$ [as in Fig. 7(c)]. The right panels in Fig. 8 show that the energy minimizes on alternating diagonal lines that weakly polarize in the field while coexisting with strongly polarized sites through the structure, similarly to Fig. 7(b).
Finally, one may wonder whether the applied magnetic field destroys the many-body spin correlations. We expect them to be robust along well defined Neel lines. To confirm this, Fig. 9 shows the link correlations for systems as in Fig. 4(a) and (b) under an applied field. Comparing these two figures, it is clear that correlations along the strong links are indeed preserved for the applied field $V_{z}=0.05$. Interestingly, correlations along the lower edge are more affected, likely due to their fewer neighbors. Notice again that the Neel lines are parallel to the lattice vector ${\boldsymbol{a}}_{2}$ in Fig. 9(b) and 4(b) as the sites in between adjacent lines are weakly correlated to their neighbors and are then strongly polarizable by the field, as seen in Fig. 7(b). These sites behave as a collection of paramagnetic (noninteracting) polarizable spins, while the sites belonging to Neel lines form a separate set of AF 1D chains. The coexistence of these two classes of sites and the dependence of their appearance and orientation with applied strain represent the main unexpected results of this work.
@@ IV Conclusions
We have studied a kagome nanoribbon using a Hubbard model to describe how anisotropies induced by uniaxial strain break the symmetry and affect the correlations due to electronic repulsion. We find that even for relatively weak interactions, the competing correlations and frustration develop coexisting systems of strongly antiferromagnetic correlations between neighboring sites and weakly connected sites that are easily polarizable in a magnetic field. The spatial distribution of the strong links and interspersed paramagnetic sites depends strongly on the orientation of the uniaxial strain, and opens the possibility of exploring an experimentally tunable probe that can give rise to different correlated behavior in a given lattice. The need to consider charge fluctuations to define a function that quantifies correlations, especially for weak $U$ values, and that complements $f^{G}$, remains an interesting theoretical question. With the growing interest in metallic kagome materials displaying strongly correlated behavior, it may be possible to apply external strain fields and explore some of the phenomena our models describe. It would be interesting, moreover, to explore how distinct Fermi levels and proximity to van Hove singularities in different materials affect this behavior. As a final reflection, there might be a connection between our results and the spin Jahn-Teller effect. The idea that spontaneous distortions could play a role in kagome materials is appealing and deserves future investigation.
@@ Appendix A Strained couplings
Now, by defining $\alpha_{\pm}\equiv\epsilon\left(1\pm\nu\right)/2$, we can recast Eq. (7) as
@@
Appendix B Coupling ratios
Figure 10: $r(\theta)$ vs $\theta$ for $\epsilon=0.30$. Note that this quantity is U-independent. Dashed lines indicate $\theta=\pi/6,\,\pi/3$, for a minimum and maximum in $E_{0}$.
For each angle $\theta$, we display the ratio $r(\theta)\equiv J_{\text{min}}(\theta)/J_{\text{max}}(\theta)$ between the smallest and largest couplings in Fig. 10. This is a non-analytical function, with cusps located at the stationary points in Fig. 3. Consistent with Eqs. (5) and (6) the plot reflects that $r(\theta)$ values closer to one indicate greater homogeneity of couplings, thus implying higher frustration–hence the peak at, e.g., $\theta=\pi/3$ being taller than the one at $\theta=\pi/6$. We conclude that $r(\theta)$ is a much coarser way to infer frustration, and therefore not as useful indicator as those discussed in the main text.